OpenStudy (anonymous):

A chocolate chip cookie recipe asks for two and one quarter times as much flour as chocolate chips. If three and three quarters cups of flour is used, what quantity of chocolate chips would then be needed, according to the recipe?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (lilg132):

5/3?

jimthompson5910 (jim_thompson5910):

2 and one quarter = 2 + 1/4 = 2.25 3 and 3 quarters = 3 + 3/4 = 3.75 F = 2.25C 3.75 = 2.25C 3.75/2.25 = C 1.6667 = C C = 1.6667 C = 1 and 2/3